OpenStudy (anonymous):

can anyone help me with the following: (11-z)z...thank you!

OpenStudy (anonymous):

using (a+b)*c=a*c+b*c (11-z)z=11z-z^2 :)

OpenStudy (anonymous):

I was making more difficult. I had 11z^2 - 2z... Thank you!

OpenStudy (anonymous):

but I see why I had it wrong...

OpenStudy (anonymous):

Distributive Law